OpenStudy (anonymous):

express in simplest form : cubed root of a^6b^9/-64 show work :)

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{\frac{a^6b^9}{-64}}\to \frac{a^2b^3}{-4}\]